OpenStudy (anonymous):

Part a.) Find the slope of the line that contains the points, (1, 2) and (2, 5). Part b.) Using the slope that you found in part a and one of the given points, write the equation of the line. Please write your final answer in slope-intercept form.

OpenStudy (anonymous):

i got it :) i think

OpenStudy (anonymous):

y= mx+ c is the slope intecept form of a line where m is the slope. The relevant equation here is y- y1/ x- x1 = y2- y1/ x2-x1 Reducing the equation y-2/ x-1 = 2-1/ 5-2 will give you the answer